# Local wind speed forecasting at short time horizons relying on both Numerical Weather Prediction and observations from surrounding stations

### v1

@@ Abstract
This study presents a hybrid neural network model for short-term (1-6 hours ahead) surface wind speed forecasting, combining Numerical Weather Prediction (NWP) with observational data from ground weather stations. It relies on the MeteoNet dataset, which includes data from global (ARPEGE) and regional (AROME) NWP models of the French weather service and meteorological observations from ground stations in the French Mediterranean. The proposed neural network architecture integrates recent past station observations (over last few hours) and AROME and ARPEGE predictions on a small subgrid around the target location. The model is designed to provide both deterministic and probabilistic forecasts, with the latter predicting the parameters of a suitable probability distribution that notably allows us to capture extreme wind events. Our results demonstrate that the hybrid model significantly outperforms baseline methods, including raw NWP predictions, persistence models, and linear regression, across all forecast horizons. For instance, the model reduces RMSE by up 30% compared to AROME predictions. Probabilistic forecasting further enhances performance, particularly for extreme quantiles, by estimating conditional quantiles rather than relying solely on the conditional mean. Fine-tuning the model for specific stations, such as those in the Mediterranean island of Corsica, further improves forecasting accuracy. Our study highlights the importance of integrating multiple data sources and probabilistic approaches to improve short-term wind speed forecasting. It defines an effective approach, even in a complex terrain like Corsica where localized wind variations are significant.
@@ 1 Introduction
Wind speed forecasting is an important issue that concerns many applications, from public safety to renewable energy production. In the energy sector, where the production of wind farms is inherently variable, accurate wind forecasts over short-term horizons (few hours) are crucial for enabling grid operators and wind farm managers to balance supply and demand, plan operations effectively, and mitigate risks (Shaw et al., 2019). Similarly, reliable short-term wind speed forecasts are essential for protecting public welfare by providing critical information for aviation, shipping, and other weather-sensitive human activities, especially in the case of severe wind episodes (Pinto et al., 2019). However, the turbulent and intermittent nature of wind, combined with rapid variations in local conditions, makes short-term wind prediction a highly complex task (Pantillon et al., 2020).
Traditional weather forecasting relies heavily on Numerical Weather Prediction (NWP) models, which simulate atmospheric dynamics based on physical principles and use initial conditions derived from observational data.
@@
National weather services usually combine NWP models at the global and regional scales to produce forecasts. For instance, Météo France uses the ARPEGE and AROME models: ARPEGE (Courtier et al., 1994), a global model with a refined resolution of about 5 km over Europe and France, provides broader-scale forecasts up to 5 days ahead but is limited in its ability to capture local phenomena; AROME (Seity et al., 2011) is a high-resolution mesoscale model downscaled from ARPEGE, specifically designed for shorter-range forecasts up to 2 days ahead and with a horizontal resolution of 1.3 km, which makes it effective in resolving weather phenomena at smaller scales and deep convection in particular. Despite the considerable progress made in recent decades (Bauer et al., 2015), NWPs are computationally intensive and often struggle to provide accurate predictions at fine spatial and temporal scales. Designed primarily for meso to synoptic scales $\mathcal{O}$(10–1000 km), NWP models are less effective in capturing localized variations and rapidly evolving phenomena such as wind gusts (Pantillon et al., 2018) or local convective episodes (Yano et al., 2018). Moreover, since the atmosphere is a chaotic system, it is desirable to provide forecasts with an estimate of the associated uncertainty, that is, to make probabilistic predictions. For this reason, operational centers produce ensemble forecasts, which consist of multiple runs of the same NWP, obtained using stochastically perturbed initial conditions and physical equations, or by aggregating results from different models (or both; Swinbank et al., 2016), a fact that makes NWP forecasts even more costly.
@@
To address those limitations of NWP models for wind speed “nowcasting” (namely, forecasting at horizons up to few hours) a range of statistical, Machine Learning (ML) and Deep Learning (DL) methods has emerged. These methods mainly rely on statistical inference or time-series model based approaches that use historical data and focus on past observed patterns to forecast future observations (see e.g., Tascikaraoglu and Uzunoglu (2014) and references therein). Some of these techniques involve the development of specific stochastic models in order to account for observed fluctuations, particularly within the realm of time series analysis (Baïle et al., 2011). Most of the recent approaches leverage more advanced ML and DL techniques. In particular, artificial Neural Networks (NN) have proven to be very efficient in capturing the nonlinear dependencies that define wind dynamics. Numerous neural network models with many different architectures have been proposed so far. The most commonly used models are Recurrent Neural Networks, temporal or spatial Convolutional Networks, Graph Convolutional Networks or transformer based models. A comprehensive overview of this rapidly evolving field can be found in numerous review papers (see e.g., Hanifi et al. (2020); Wang et al. (2021)).
@@
Besides pure NWP or pure statistical approaches, hybrid methods that combine the two have been experimented in order to produce better predictions, especially at short-term (Huang et al., 2012). Hybrid models, as the name suggests, aim to take advantage of the strengths of both NWP and statistical methods. These models may incorporate NWP output as predictors within statistical frameworks, such as time series models or ML approaches (like e.g., NN models). In these cases, one expects that the NWP forecasts provide valuable information about large-scale atmospheric conditions, while processing observational data allows one to account for local effects, model biases, and unresolved sub-grid processes to improve wind speed or wind power nowcasting (Hoolohan et al., 2018; Brabec et al., 2021; Donadio et al., 2021; He et al., 2022). Let us notice that such approaches are very close to the so-called statistical post-processing of NWP forecasts (Vannitsem et al., 2021). Indeed, it is well known that direct forecasting from NWP, regardless of whether it is global or regional, often exhibits systematic biases and limitations, particularly at the local scale and close to the surface for variables like wind speed (Zamo et al., 2016). These errors stem from factors such as insufficient model resolution, incomplete representation of physical processes, and inaccuracies in initial conditions. For that reason, operational centers consider various post-processing methods in order to remove those biases and improve prediction. The most commonly used method is the “Model Output Statistics” (MOS) which consists of regressing observed data from raw NWP outputs. This framework has been extended to probabilistic forecasting with the “Ensemble Model Output Statistic” (EMOS) that was designed to handle ensemble predictions (Gneiting et al., 2005). More elaborated statistical methods, that can capture non-linear, complex relationships between NWP outputs and observed weather variables like Kalman filtering (Pelosi et al., 2017), random forests (Zamo et al., 2016) or NN models (Rasp and Lerch, 2018; Veldkamp et al., 2021) have been proposed more recently. Consequently, advancements in NWP post-processing methods and hybrid model approaches have made them increasingly difficult to distinguish, as both allow for improved short-term forecasting accuracy by integrating NWPs with local observational data.
@@
This study contributes to the development of hybrid statistical methods for reliable nowcasting of surface wind speed. It leverages the MeteoNet database, an open-access resource provided by Météo France, which includes both NWP and observational data (Larvor et al., 2020). MeteoNet covers two regions of France—the Southeast and Northwest—and provides essential meteorological variables at hourly and sub-hourly temporal resolutions, making it well-suited for training and validating ML models for short-term, location-specific forecasts. We focus on the French Mediterranean region, particularly Corsica, a French Mediterranean island (80 × 180 km) characterized by complex mountainous terrain and frequent high-impact meteorological events such as windstorms, heavy rainfall, thunderstorms and Saharan dust incursions (Scheffknecht et al., 2017; Coquillat et al., 2019; Lfarh et al., 2023). The Mediterranean region is also a recognized climate change hotspot, motivating various research initiatives (e.g., Drobinski et al., 2014; Hatzaki et al., 2023). We introduce a NN model to enhance short-term wind speed forecasting in this geographically complex setting. Our model integrates NWP data from both global (ARPEGE) and regional (AROME) models with station observations to generate deterministic and probabilistic forecasts for lead times of one to six hours at any given selected station. By combining the strengths of physically based NWP models and data-driven ML techniques, our proposed framework aims at capturing both large-scale weather patterns and localized wind variations in order to improve forecasting accuracy and adaptability to site-specific conditions. Our framework is designed to support operational needs in the wind energy sector, providing actionable insights for planning, risk mitigation, and decision-making.
@@
Let us end this introductory section by mentioning that a similar approach, utilizing the same dataset, was recently proposed by Marcille et al. (2024) in order to forecast wind speed at offshore locations where observational data is unavailable.
The paper is organized as follows. In Section 2 we describe the MeteoNet dataset used in this work. Section 3 is devoted to the description of the forecasting task we focus on and of various models which performance are compared. We notably introduce the ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$ models that leverage both NWP and station MeteoNet data for producing respectively deterministic and probabilistic short-term surface wind speed forecasting. All the results are provided in Section 4. Subsection 4-4.1 presents a comparative performance analysis of ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$ against various baselines for all considered stations, while 4-4.2 focuses on the results of versions of the model specifically fine-tuned for sites in Corsica. Section 5 summarizes our findings and proposes various prospects for future research.
@@ 2.1 The MeteoNet database
The results presented in this article are based on the MeteoNet database, which is a comprehensive and freely available meteorological dataset published by Météo France, the French national meteorological service (Larvor et al., 2020). It has been specifically designed to support research and development in meteorology, climatology, and machine learning applications. The MeteoNet database contains high-resolution meteorological data over, respectively, the Northwest and Southeast quarters of the France territory covering the period from 01-01-2016 to 12-31-2018. It includes observations from ground-based weather stations, gridded forecasts from NWP models, and auxiliary geospatial information.
@@
In this work, we will consider exclusively the Southeast data, influenced mostly by a Mediterranean climate with, as mentioned in the introduction, a special focus on sites spread all over Corsica. Observation data are provided by the Météo France ground station network, with a large number of stations (278 stations with a fraction of available data above a given threshold). The position of each of these stations is shown in Fig. 1(a). They provide, at a time resolution of 6 min, times records of different weather variables as notably 10 $m$ wind speed and direction and 2 $m$ air temperature. In this study, we also use the forecasts data from Météo France operational NWP models ARPEGE and AROME included in the MeteoNet database with 0.1 and 0.025 horizontal resolutions, respectively. A single run is provided for each day, the one beginning at 00 UTC that spans the forecast ranges from 00 to 24 UTC, with a time step of 1 h for the AROME data and 1 h or 3 h for the ARPEGE data, depending on the time horizon. AROME data are composed of different surface fields (notably $10$ m wind speed, $2$ m temperature, relative humidity and mean-sea-level pressure) whereas ARPEGE provides predictions of various fields (like horizontal wind speed components or temperature) at 7 different isobar levels spanning the low to mid troposphere and at 7 different heights in the lower troposphere for the pressure variable. The precise way we use a subset of these data in our model is described in Sec. 3.
@@ 2.2 Wind speed station data in Corsica
For comparison purpose, we also utilize a specific dataset of wind speed and direction time series previously used in Baggio and Muzy (2024). This dataset comprises hourly wind speed measurements recorded by Météo France at various locations across Corsica from January 1, 2011, to December 31, 2020. In the following, we focus on four specific sites—Ajaccio, Lucciana, Renno, and Figari—to compare the proposed approach with previous machine learning methods, such as the one introduced in Baïle and Muzy (2023), which relies exclusively on past observed data from the target station and its 15 closest neighboring stations. These four selected sites together with their neighboring stations have a low proportion of missing data, with less than 15% of the total possible data points unavailable. For this analysis, the period from January 1, 2011, to December 31, 2015, is used as the training sample, while the comparison is performed on the validation and test subsets from the MeteoNet 2016–2018 period.
3 Forecasting task and models
Our goal in this paper is to propose a model that provides wind speed forecasts at some single site over short term lead time of few hours. In this section, we set the main notations and describe this specific prediction goal in mathematical terms. We then introduce the main model designed to tackle this problem, along with alternative baseline models used for comparison.
3.1 Prediction at multiple horizons

### v2

# Local wind speed forecasting at short time horizons based on Numerical Weather Prediction and observations from surrounding stations

@@ Abstract
This study presents a hybrid neural network model for short-term (1-6 hours ahead) surface wind speed forecasting, combining Numerical Weather Prediction (NWP) with observational data from ground weather stations. It relies on the MeteoNet dataset, which includes data from global (ARPEGE) and regional (AROME) NWP models of the French weather service and meteorological observations from ground stations in the French Mediterranean. The proposed neural network architecture integrates station observations from the last few hours and AROME and ARPEGE predictions on a small subgrid around the target location. The model is designed to provide both deterministic and probabilistic forecasts, with the latter predicting the parameters of a suitable probability distribution that notably allows us to capture extreme wind events. Our results demonstrate that the hybrid model significantly outperforms baseline methods, including raw NWP predictions, persistence models, and linear regression, across all forecast horizons. For instance, the model reduces RMSE by up to 30% compared to AROME predictions. Probabilistic forecasting further enhances performance, particularly for extreme quantiles, by estimating conditional quantiles rather than relying solely on the conditional mean. Fine-tuning the model for specific stations, such as those in the Mediterranean island of Corsica, further improves forecasting accuracy. Our study highlights the importance of integrating multiple data sources and probabilistic approaches to improve short-term wind speed forecasting. It defines an effective approach, even in a complex terrain like Corsica where localized wind variations are significant.
@@ 1 Introduction
Wind speed forecasting is an important issue that concerns many applications, from public safety to renewable energy production. In the energy sector, where the production of wind farms is inherently variable, accurate wind forecasts over short-term horizons (few hours) are crucial for enabling grid operators and wind farm managers to balance supply and demand, plan operations effectively, and mitigate risks (Shaw et al., 2019). Likewise, accurate short-term wind speed forecasts are critical for safeguarding public welfare by supplying essential information for aviation, maritime operations, and other weather-sensitive activities, especially in the case of severe wind episodes (Pinto et al., 2019). However, the turbulent and intermittent nature of wind, combined with rapid variations in local conditions, makes short-term wind prediction a highly complex task (Pantillon et al., 2020).
@@
Traditional weather forecasting relies heavily on Numerical Weather Prediction (NWP) models, which simulate atmospheric dynamics based on physical principles and use initial conditions derived from observational data. National weather services usually combine NWP models at the global and regional scales to produce forecasts. For instance, Météo France uses the ARPEGE and AROME models: ARPEGE (Courtier et al., 1994), a global model with a refined resolution of about 5 km over Europe and France, provides broader-scale forecasts up to 5 days ahead but is limited in its ability to capture local phenomena; AROME (Seity et al., 2011) is a high-resolution limited-area model downscaled from ARPEGE, specifically designed for shorter-range forecasts up to 2 days ahead and with a horizontal resolution of 1.3 km, which makes it effective in resolving weather phenomena at smaller scales and deep convection in particular. Despite the considerable progress made in recent decades (Bauer et al., 2015), NWPs are computationally intensive and often struggle to provide accurate predictions at fine spatial and temporal scales. Designed primarily for meso to synoptic scales $\mathcal{O}$(10–1000 km), NWP models are less effective in capturing localized variations and rapidly evolving phenomena such as wind gusts (Pantillon et al., 2018) or local convective episodes (Yano et al., 2018). Moreover, since the atmosphere is a chaotic system, it is desirable to provide forecasts with an estimate of the associated uncertainty, that is, to make probabilistic predictions. For this reason, operational centers produce ensemble forecasts, which consist of multiple runs of the same NWP, obtained using stochastically perturbed initial conditions and physical equations, or by aggregating results from different models (or both; Swinbank et al. (2016)), a fact that makes NWP forecasts even more costly.
@@
To address those limitations of NWP models for wind speed “nowcasting” (namely, forecasting at horizons up to few hours) a range of statistical, Machine Learning (ML) and Deep Learning (DL) methods has emerged. These methods mainly rely on statistical inference or time-series model based approaches that use historical data and focus on past observed patterns to forecast future observations (see, e.g., Tascikaraoglu and Uzunoglu (2014) and references therein). Some of these techniques involve the development of specific stochastic models in order to account for observed fluctuations, particularly within the realm of time series analysis (Baïle et al., 2011). Most of the recent approaches leverage more advanced ML and DL techniques. In particular, artificial Neural Networks (NN) have proven to be very efficient in capturing the nonlinear dependencies that define wind dynamics. Numerous neural network models with many different architectures have been proposed so far. The most commonly used models are Recurrent Neural Networks, temporal or spatial Convolutional Networks, Graph Convolutional Networks or transformer based models. A comprehensive overview of this rapidly evolving field can be found in numerous review papers (see, e.g., Hanifi et al. (2020); Wang et al. (2021)).
Besides pure NWP or pure statistical approaches, hybrid methods that combine the two have been experimented in order to produce better predictions, especially at short-term (Huang et al., 2012). In this paper, we define hybrid models as machine learning-based architectures trained on both NWP forecasts and other source of data (like observational data), where the model structure explicitly allows for the fusion of these inputs—potentially through separate branches or feature encoders. These models may incorporate NWP output as predictors within statistical frameworks, such as time series models or ML approaches (like e.g., NN models). One expects that the NWP forecasts provide valuable information about large-scale atmospheric conditions, while processing observational data allows one to account for local effects, model biases, and unresolved sub-grid processes to improve wind speed or wind power nowcasting (Hoolohan et al., 2018; Brabec et al., 2021; Donadio et al., 2021; He et al., 2022).
@@
It is noteworthy that such approaches should not be confused with the statistical post-processing of NWP forecasts (Vannitsem et al., 2021). Indeed, it is well known that direct forecasting from NWP, regardless of whether it is global or regional, often exhibits systematic biases and limitations, particularly at the local scale and close to the surface for variables like wind speed (Zamo et al., 2016). These errors stem from factors such as insufficient model resolution, incomplete representation of physical processes, and inaccuracies in initial conditions. For that reason, operational centers consider various post-processing methods in order to remove those biases and improve prediction. The most commonly used method is the “Model Output Statistics” (MOS), which consists of regressing observed data from raw NWP outputs. This framework has been extended to probabilistic forecasting with the “Ensemble Model Output Statistic” (EMOS), designed to handle ensemble predictions (Gneiting et al., 2005). However, MOS and EMOS typically rely on linear statistical corrections and do not incorporate other data, e.g., observational time series, as separate input features. More elaborated statistical methods, that can capture non-linear, complex relationships between NWP outputs and observed weather variables—like Kalman filtering (Pelosi et al., 2017), random forests (Zamo et al., 2016), or NN models (Rasp and Lerch, 2018; Veldkamp et al., 2021)—have been proposed more recently. Consequently, advancements in NWP post-processing methods and hybrid model approaches have made them increasingly difficult to distinguish. Still, the key distinction we adopt in this work is that hybrid models aim to jointly learn from various types of forecasts and observations within an integrated learning framework, whereas post-processing focuses on correcting NWP forecasts—often using simpler statistical models.
@@
This study contributes to the development of hybrid statistical methods for reliable nowcasting of surface wind speed. It leverages the MeteoNet database, an open-access resource provided by Météo France, which includes both NWP and observational data (Larvor et al., 2025). MeteoNet covers two regions of France (the Southeast and Northwest) and provides essential meteorological variables at hourly and sub-hourly temporal resolutions, making it well-suited for training and validating ML models for short-term, location-specific forecasts. We focus on the French Mediterranean region, particularly Corsica, a French Mediterranean island (80 × 180 km) characterized by complex mountainous terrain and frequent high-impact meteorological events such as windstorms, heavy rainfall, thunderstorms and Saharan dust incursions (Scheffknecht et al., 2017; Coquillat et al., 2019; Lfarh et al., 2023). The Mediterranean region is also a recognized climate change hotspot, motivating various research initiatives (e.g., Drobinski et al. (2014), Hatzaki et al. (2023)). We introduce a NN model to enhance short-term wind speed forecasting in this geographically complex setting. Our model integrates NWP data from both global (ARPEGE) and regional (AROME) models with station observations to generate deterministic and probabilistic forecasts for lead times of one to six hours at any given selected station. By combining the strengths of physically based NWP models and data-driven ML techniques, our proposed framework aims at capturing both large-scale weather patterns and localized wind variations in order to improve forecasting accuracy and adaptability to site-specific conditions. Our framework is designed to support operational needs in the wind energy sector, providing actionable insights for planning, risk mitigation, and decision-making.
@@
Finally, it is relevant to mention that a similar approach utilizing the same dataset was recently proposed by Marcille et al. (2024) to forecast wind speed at offshore locations where observational data is unavailable.
The paper is organized as follows. Section 2 introduces the MeteoNet dataset used throughout this study. Section 3 defines the forecasting task under consideration, along with the associated evaluation framework. Section 4 presents the different forecasting models and baselines included in our comparison. In particular, we introduce two hybrid architectures, denoted by $\mathcal{M}$ and $\mathcal{M}^{\prime}$, which combine NWP outputs with weather station data from MeteoNet to produce deterministic and probabilistic short-term forecasts of near-surface wind speed, respectively. Section 5 reports the empirical results. Subsection 5.1 provides a comparative performance analysis of $\mathcal{M}$ and $\mathcal{M}^{\prime}$ against several baselines across all stations. Subsection 5.2 focuses on models specifically fine-tuned for stations located in Corsica, where we evaluate both standard point forecasting (based on error minimization) and the ability to predict extreme wind conditions through threshold exceedance events. Section 6 presents limitations of the proposed approach, identifies potential improvements and considers some operational aspects. Finally, Section 7 summarizes our main findings and suggests directions for future work.
@@ 2.1 The MeteoNet database
The results presented in this article are based on the MeteoNet database, which is a comprehensive and freely available meteorological dataset published by Météo France, the French national meteorological service (Larvor et al., 2025). It has been specifically designed to support research and development in meteorology, climatology, and machine learning applications. The MeteoNet database contains high-resolution meteorological data over, respectively, the Northwest and Southeast quarters of the France territory covering the period from 1 January 2016 to 31 December 2018. It includes observations from ground-based weather stations, gridded forecasts from NWP models, and auxiliary geospatial information.
@@
In this work, we will consider exclusively the Southeast data, influenced mostly by a Mediterranean climate with, as mentioned in the introduction, a special focus on sites spread all over Corsica. Observation data are provided by the Météo France ground station network. A subset of 278 stations was selected by setting a data availability threshold to ensure statistical significance. The position of each of these selected stations is shown in Fig. 1(a). They provide, at a time resolution of 6 min, time records of different weather variables as notably 10 $m$ wind speed and direction and 2 $m$ air temperature. In this study, we also use the forecast data from Météo France operational NWP models ARPEGE and AROME included in the MeteoNet database with 0.1 and 0.025 horizontal resolutions, respectively. A single run is provided for each day, the one beginning at 00 UTC that spans the forecast ranges from 00 to 24 UTC, with a time step of 1 h for the AROME data and 1 h or 3 h for the ARPEGE data, depending on the time horizon. AROME data are composed of different surface fields (notably $10$ m wind speed, $2$ m temperature, relative humidity and mean-sea-level pressure) whereas ARPEGE provides predictions of various fields (like horizontal wind speed components or temperature) at 7 different isobar levels spanning the low to mid troposphere and at 7 different heights in the lower troposphere for the pressure variable.
To prepare the dataset for use in the training, validation, and test phases of the models presented in Sec. 4, we implement a two-step data preprocessing workflow. In the first step, the raw MeteoNet Southeast dataset is reorganized into a station-centric format. Specifically, for each available station and each day, we generate a NetCDF file that compiles all relevant data: the station’s ground-based time series for the day (including context from adjacent days), as well as AROME and ARPEGE model outputs in the form of spatial sub-images centered on the station. This stage is purely structural, and no filtering or cleaning is applied, in order to preserve the integrity of the original data. The second step, which is closely tied to the specific prediction task defined in Sec. 3, involves transforming these NetCDF files into model-ready inputs. Each file is read and processed to construct the input tensors and corresponding output values used during training, validation, and testing (see Sec. 4.1 for details on the input structure). During this stage, we discard any samples (i.e., specific station-day pairs) containing missing or invalid values, normalize the input features, and store the resulting data in a unified format—either as a large binary file or an in-memory arrays —compatible with TensorFlow or PyTorch data generators. This two-stage process ensures flexibility, reproducibility, and robust quality control across the entire data pipeline.
@@ 2.2 Wind speed station data in Corsica
For comparison purpose, we also utilize a specific dataset of wind speed and direction time series previously used in Baggio and Muzy (2024). This dataset comprises hourly wind speed measurements recorded by Météo France at various locations across Corsica from 1 January 2011 to 31 December 2020. It can be used to train a neural network model such as the one introduced in Baïle and Muzy (2023) which relies exclusively on past observed data from the target station and its 15 closest neighboring stations. In section 5.2.1, we compare a model relying exclusively on observed data to the hybrid approach advocated in this paper on four representative sites in Corsica, namely Ajaccio, Lucciana, Renno, and Figari. These four selected sites together with their neighboring stations have a low proportion of missing data, with less than 15% of the total possible data points unavailable. For this analysis, the period from 1 January 2011 to 31 December 2015 is used as the training sample for model of Baïle and Muzy (2023), while the comparison is performed on the same validation and test subsets we define for the MeteoNet dataset over the 2016–2018 period (see Sec. 4.4).
3 Prediction task at multiple horizons
In this section, we establish the primary notations and formally define the forecasting problem addressed in this study. Our main objective is to predict the surface wind speed at a specific site over short-term forecast horizons, typically ranging from 1 to 6 hours ahead. As explained in the introduction, a central motivation for this task, beyond producing accurate wind speed forecasts, is to enable the assessment of specific events such as whether the wind speed will exceed a predefined extreme value threshold—information that is critical for decision-making in weather-sensitive applications. To model this predictive task, we consider two alternative strategies. The first is a deterministic approach, which yields directly a point forecast—i.e., a single predicted value—for future wind speed. The second is a probabilistic approach, which estimates the full conditional distribution of future wind speeds, thereby supporting both point predictions and probabilistic quantification of events such as threshold exceedance. We evaluate and compare the performance of these approaches using appropriate metrics.